# Finding Most Shattering Minimum Vertex Cuts of Polylogarithmic Size in Near-Linear Time

### v1

@@ 5.4 Additional Lemmas
Suppose that Algorithm˜1 on input $(x,\nu,\Pi)$ returns an unverified set $U$ such that $U\neq\varnothing$. Then, $U$ is a $k$-separator. Specifically, let $z$ be the far-most endpoint of an arbitrary path $\pi\in\Pi$. Then, $x$ and $z$ are not connected in $G\setminus U$.
@@
Suppose that Algorithm˜1 returns a nonempty unverified set $U$ on input $(x,\nu,\Pi)$. Let $Q_{x}$ denote the component of $G\setminus U$ containing $x$. There exists a modified version of Algorithm˜1 that also computes $\mathrm{vol}(Q_{x})$. The modification requires $\mathcal{O}(k\nu)$ additional time, subsumed by the runtime of Algorithm˜1.
@@ 6 Resolving Unverified Sets
We now address the case where Algorithm˜1 returns an unverified set $U$ on input $(x,\nu,\Pi)$. The difference between $U$ and the list of returned $k$-shredders $\mathcal{L}$ is that the algorithm did not find a component $Q$ of $G\setminus\Pi$ such that $N(Q)=U$. In some sense, this means that we “lose” a component of $G\setminus U$. The danger in guessing that $U$ is a $k$-shredder is that $U$ might only be a $k$-separator. Imagine that $G\setminus U$ has exactly two components: $Q_{1}$ and $Q_{2}$. If $\mathrm{vol}(Q_{2})\gg\mathrm{vol}(Q_{1})$, then it becomes quite challenging to determine whether there exists a third component of $G\setminus U$ in $\mathcal{O}(\mathrm{vol}(Q_{1}))$ time. To cope with this uncertainty, we make another structural classification.
@@
One aspect of this definition may seem strange: we have specifically described $\nu$ as a power of two. This is because in the later sections, we will use geometric sampling to capture $k$-shredders. The sampling parameters we use will be powers of two. Hence, we have imposed this slightly arbitrary detail on Definition˜6.1. For now, all that matters is that $\frac{1}{2}\nu<\mathrm{vol}(\mathcal{R})\leq\nu$.
@@
There are two main lemmas for this section. Lemma˜6.3 handles low-degree unverified sets and is presented in Section˜6.1. Lemma˜6.9 handles high-degree unverified sets and is presented in Section˜6.2. The idea is that after Algorithm˜1 returns an unverified set $U$, we will use Lemma˜6.3 to test whether $U$ is a low-degree $k$-shredder. If not, we will leave keep $U$ as a potential high-degree $k$-shredder. After all unverified sets have been returned, we use Algorithm˜3 to extract all high-degree $k$-shredders from the remaining unverified sets.
@@ 6.1 Low-Degree
Suppose that $U$ is a low-degree $k$-shredder with partition $(C,\mathcal{R})$. Suppose that the tuple $(x,\nu,\Pi)$ captures $U$ and Algorithm˜1 reports $U$ as an unverified set. Our goal is to design an algorithm that confirms $U$ is a $k$-shredder in the same time complexity as Algorithm˜1 up to $\text{polylog}(n)$ and $\text{poly}(k)$ factors.
Lemma˜5.11 gives us two vertices in different components of $G\setminus U$: $x$ and $z$ (where $z$ is the far-most endpoint of any path in $\Pi$). Since $U$ is a $k$-shredder, there must exist a third component of $G\setminus U$ and every vertex in $U$ must be adjacent to a vertex in this third component. Because $U$ is low-degree, there must exist a vertex $u\in U$ with $\deg(u)\leq\nu$. The idea is to scan through the edges adjacent to this low-degree vertex $u$ and find an edge $(u,y)$ such that $y$ is neither connected to $x$ nor $z$ in $G\setminus U$. To make the scanning procedure viable, we need to efficiently answer pairwise connectivity queries in $G\setminus U$. The key ingredient is a result obtained by Kosinas in [Kos23]. Specifically, the following holds.
@@
We defer the method of capturing $k$-shredders for later. For now, let us assume that the tuple $(x,\nu,\Pi)$ captures a $k$-shredder $U$ and $U$ is reported as an unverified set by Algorithm˜1. We show an auxiliary algorithm that will be helpful in listing low-degree $k$-shredders.
@@
After preprocessing the input graph using the data structure from ˜6.2, there exists a deterministic algorithm that takes in as input $(x,\nu,\Pi,U)$, where $U$ is an unverified set returned by Algorithm˜1 on input $(x,\nu,\Pi)$. The algorithm outputs true if $U$ is a $k$-shredder and there exists a vertex $u\in U$ such that $\deg(u)\leq\nu$ in $\mathcal{O}(k^{4}\log n+k\nu)$ time.
@@
In [LS22], Long and Saranurak showed the existence of a pairwise connectivity oracle subject to vertex failures with $\mathcal{O}(k^{2}n^{o(1)})$ update time and $\mathcal{O}(k)$ query time. Our usage of Kosinas’s oracle involves making one update operation with a set of size $k$ and $\mathcal{O}(\nu)$ many connectivity queries. Using Long and Saranurak’s version, we can improve the dependency on $k$ in the time complexity of Algorithm˜2 to $\mathcal{O}(k^{2}n^{o(1)}+k\nu)$ time.
@@
As mentioned above, the idea is to scan through the edges adjacent to $u$ in order to find a vertex $y$ such that $y$ is neither connected to $x$ nor $z$ in $G\setminus U$ (see Figure˜6). We will be utilizing ˜6.2 to determine whether a pair of vertices $(x,y)$ are connected in $G\setminus U$. Pseudocode is given in Algorithm˜2.
@@
1: $x$ - a sample vertex $\nu$ - volume parameter $\Pi$ - a set of $k$ openly-disjoint paths starting from $x$ $U$ - unverified set returned by Algorithm˜1 on $(x,\nu,\Pi)$
2: true if (1) and (2) are satisfied, false otherwise $Z$ (1) $U$ is a $k$-shredder $Z$ (2) there exists a vertex $u\in U$ such that $\deg(u)\leq\nu$
3: if all vertices in $U$ have degree greater than $\nu$ then
4:   return false
5: $u\leftarrow$ a vertex in $U$ with $\deg(u)\leq\nu$
6: $\pi\leftarrow$ the path in $\Pi$ containing $u$
7:
$z\leftarrow$ the far-most endpoint of $\pi$
8:
9: $\Phi\leftarrow$ the pairwise connectivity oracle from ˜6.2 (already preprocessed)
10: update $\Phi$ on vertex failure set $U$
11: for each edge $(u,y)$ adjacent to $u$ such that $y\notin U$ do
12:   $q_{x,y}\leftarrow$ connectivity query between $x$ and $y$ in $G\setminus U$
13:   $q_{y,z}\leftarrow$ connectivity query between $y$ and $z$ in $G\setminus U$
14:   if $q_{x,y}$ is false and $q_{y,z}$ is false then
15:    return true
@@
Let $(x,\nu,\Pi,U)$ be an input tuple to Algorithm˜2. Then, the algorithm returns true if and only if $U$ is a $k$-shredder and there exists a vertex $u\in U$ such that $\deg(u)\leq\nu$.
@@
Suppose that $U$ is a $k$-shredder and that there exists a vertex $u\in U$ such that $\deg(u)\leq\nu$. Let $u$ be the vertex found by line 5. Let $\pi$ be the unique path in $\Pi$ that contains $u$ and let $z$ be the far-most endpoint of $\pi$. These are found by lines 6-7. By Lemma˜5.11, we have at least two components of $G\setminus U$ so far: $Q_{x}$ (the component containing $x$) and $Q_{z}$ (the component containing $z$). Since $U$ is a $k$-shredder, there exists a third component $Q_{y}$ of $G\setminus U$. Let $y$ be a vertex in $Q_{y}$ adjacent to $u$. Note that $x,y,z$ are each in distinct components of $G\setminus U$, which implies they are pairwise disconnected in $G\setminus U$. We will find $(u,y)$ by scanning through the edges adjacent to $u$. Prior to calling the algorithm, we will initialize a global pairwise connectivity oracle subject to vertex failures as in ˜6.2. We update the oracle with respect to $U$ on line 10 so we can answer pairwise connectivity queries in $G\setminus U$. The rest is straightforward. The for loop on line 11 iterates through all edges adjacent to $u$. Hence, the edge $(u,y)$ described above will be found. For each edge $(u,y)$, we perform the two pairwise connectivity queries in $G\setminus U$ by lines 12-13. Lastly, we will discover that $x,y,z$ are all pairwise disconnected in $G\setminus U$ and confirm that $U$ is a $k$-shredder by line 14.
Now suppose that $U$ is not a $k$-shredder or there does not exist a vertex $u\in U$ with $\deg(u)\leq\nu$. If $U$ is not a $k$-shredder, then there are exactly two components in $G\setminus U$: the component containing $x$, and the component containing $z$ (the far-most endpoint of any path in $\Pi$). Therefore, for every edge $(u,y)$ we iterate through such that $u\in U,y\notin U$, either $x$ and $y$ are connected in $G\setminus U$ or $y$ and $z$ are connected in $G\setminus U$. Line 14 guarantees that we will not return true. If there does not exist a vertex $u\in U$ with $\deg(u)\leq\nu$, we can return false immediately. This is done by line 3. ∎
@@
Algorithm˜2 on input $(x,\nu,\Pi,U)$ runs in $\mathcal{O}(k^{4}\log n+k\nu)$ time.
@@
Line 3 can be implemented in $\mathcal{O}(k)$ time by iterating over all vertices in $U$ and checking the sizes of their adjacency lists. Lines 5-7 can be done in constant time. The update on line 10 takes $\mathcal{O}(k^{4}\log n)$ time according to ˜6.2. Iterating over the edges adjacent to $u$ takes $\mathcal{O}(\nu)$ time because $\deg(u)\leq\nu$. For each edge, we make two pairwise connectivity queries, each of which takes $\mathcal{O}(k)$ time according to ˜6.2. This gives the total time bound of $\mathcal{O}(k)+\mathcal{O}(k^{4}\log n)+\mathcal{O}(k\nu)=\mathcal{O}(k^{4}\log n+k\nu)$. ∎
Proof of Lemma˜6.3.
The correctness and time complexity of Algorithm˜2 are given by Lemma˜6.5 and Lemma˜6.6. ∎
@@ 6.2 High-Degree
Now we have a tool to obtain a vertex in $\mathcal{R}$. We are not finished yet, as we need to connect this idea with the unverified sets reported by Algorithm˜1. Suppose that $U$ was reported by Algorithm˜1 on input $(x,\nu,\Pi)$. We have that $x$ is in some component $Q$ of $G\setminus U$. With Lemma˜6.7, our goal is to sample a vertex $y$ in a different component $Q_{y}$ of $G\setminus U$. As described above, we can exploit the fact that $U$ is a high-degree $k$-shredder by exploring low-degree vertices in the neighborhood of $y$ to recover $Q_{y}$. At this point, we have essentially recovered two components of $G\setminus U$: $Q$ and $Q_{y}$. To make further progress, we present a short, intuitive lemma.
@@
There exists a randomized Monte Carlo algorithm that takes as input a list $\mathcal{U}$ of tuples of the form $(x,\nu,\Pi,U)$, where $U$ is an unverified set returned by Algorithm˜1 on $(x,\nu,\Pi)$. The algorithm returns a list of $k$-shredders $\mathcal{L}$ that satisfies the following. If $(x,\nu,\Pi,U)\in\mathcal{U}$ is a $k$-tuple such that $U$ is a high-degree $k$-shredder, then $U\in\mathcal{L}$ with probability $1-n^{-100}$. The algorithm runs in $\mathcal{O}(k^{2}m\log^{2}n)$ time.
@@
We summarize the argument made above. Fix a tuple $(x,\nu,\Pi,U)$ in $\mathcal{U}$ such that $U$ is a high-degree $k$-shredder with partition $(C,\mathcal{R})$. Let $\nu$ denote the unique power of two satisfying $\frac{1}{2}\nu<\mathrm{vol}(\mathcal{R})\leq\nu$. Because $U$ is high-degree, we have for all $u\in U$, $\deg(u)>\nu$. Let $Q_{x}$ denote the component of $G\setminus U$ containing $x$. Suppose we have obtained a vertex $y$ in a component $Q_{y}\in\mathcal{R}\setminus\{Q_{x}\}$. Since all vertices in $Q_{y}$ have degree at most $\nu$, the idea is to explore all vertices connected to $y$ that have degree at most $\nu$. Because $N(Q_{y})=U$ and each vertex in $U$ has degree greater than $\nu$, we will compute $Q_{y}$ as the set of explored vertices. After computing $Q_{y}$, we will have obtained two components of $G\setminus U$: $Q_{x}$ and $Q_{y}$. In order to confirm that $U$ is a $k$-shredder, all that is left to do is to make sure that $Q_{x}$ and $Q_{y}$ are not the only components of $G\setminus U$. Lemma˜6.8 implies this can be done by verifying that $\mathrm{vol}(Q_{x})+\mathrm{vol}(Q_{y})+|E(U,U)|<m$. Now, the final step is to find such a vertex $y$. We use a random sampling procedure for this task. Consider Algorithm˜3. The following two lemmas prove the correctness and time complexity of Algorithm˜3.
@@
1: $\mathcal{U}$ - a list of tuples of the form $(x,\nu,\Pi,U)$
2: $\mathcal{L}$ - a list of $k$-shredders that satisfies the following: $Z$ if $(x,\nu,\Pi,U)\in\mathcal{U}$ is such that $U$ is a high-degree $k$-shredder, $Z$ then $U\in\mathcal{L}$ with probability $1-n^{-100}$
3: $\mathcal{L}\leftarrow\varnothing$
4: for $i\leftarrow 0$ to $\lceil\log\left(\frac{m}{k}\right)\rceil$ do
5:   $\nu\leftarrow 2^{i}$
6:   for $400\frac{m}{\nu}\log n$ times do
7:    independently sample an edge $(y,z)$ uniformly at random
8:    $Q_{y}\leftarrow$ set of vertices explored by BFS seeded at $y$: 9:  ignore vertices $v$ such that $\deg(v)>\nu$ 10:  terminate early the moment more than $\nu$ edges are explored
11:    if BFS terminated early then
12:      skip to next sample
13:    else if there exists a tuple $(x,\nu_{x},\Pi,U)\in\mathcal{U}$ such that $N(Q_{y})=U$ then
14:      $Q_{x}\leftarrow$ component of $G\setminus U$ containing $x$
15:      if $x\notin Q_{y}$ and $\mathrm{vol}(Q_{x})+\mathrm{vol}(Q_{y})+|E(U,U)|<m$ then
16:       $\mathcal{L}\leftarrow\mathcal{L}\cup\{U\}$
17: return $\mathcal{L}$
@@
Let $\mathcal{U}$ be the list of tuples given as input to Algorithm˜3 and let $\mathcal{L}$ be the output list. Suppose $U$ is a $k$-shredder such that there exists a $k$-tuple $(x,\nu,\Pi,U)\in\mathcal{U}$. Then, $U\in\mathcal{L}$ with probability $1-n^{-100}$. Furthermore, every set in $\mathcal{L}$ is a $k$-shredder.
@@
Fix a tuple $(x,\nu,\Pi,U)$ in $\mathcal{U}$ such that $U$ is a high-degree $k$-shredder with partition $(C,\mathcal{R})$. We will show that $U\in\mathcal{L}$ with high probability. Let $Q_{x}$ denote the component in $\mathcal{R}$ containing $x$. Let $Q_{y}$ denote a component of $\mathcal{R}\setminus\{Q_{x}\}$. We will first show that the algorithm obtains a vertex $y\in Q_{y}$ with high probability. Let $\nu$ denote the unique power of two satisfying $\frac{1}{2}\nu<\mathrm{vol}(Q_{y})\leq\nu$. Lemma˜6.7 implies that if we sample $400\frac{m}{\nu}\log n$ edges, we will obtain an edge $(y,z)$ such that $y\in Q_{y}$. Although we do not know the explicit value of $\nu$, all powers of two up to $\frac{m}{\nu}$ are tried by the algorithm during the loop on line 4. Hence, we will eventually obtain an edge $(y,z)$ with $y\in Q_{y}$ with probability $1-n^{-100}$ during the for loop on line 6.
After obtaining a vertex $y\in Q_{y}$, we perform BFS seeded at $y$. Since $\frac{1}{2}\nu<\mathrm{vol}(Q_{y})\leq\nu$, we know that all vertices in $Q_{y}$ have degree at most $\nu$. We also know that $Q_{y}$ can be computed by traversing at most $\nu$ edges because $\mathrm{vol}(Q_{y})\leq\nu$. Lastly, we know that all vertices in $U$ have degree greater than $\nu$ because $U$ is a high-degree $k$-shredder. This implies $Q_{y}$ will be computed by line 8. After computing $Q_{y}$, we can determine whether there exists a tuple $(x,\nu_{x},\Pi,U)$ in the input list such that $N(Q_{y})=U$. If such a tuple exists, we should first check that $x\notin Q_{y}$. If so, we now have two components of $G\setminus U$: $Q_{x}$ and $Q_{y}$. By Lemma˜6.8, the only step left to determine whether $U$ is a $k$-shredder is verifying that $\mathrm{vol}(Q_{x})+\mathrm{vol}(Q_{y})+|E(U,U)|<m$. Since $U$ is a $k$-shredder, this inequality must be true. We can conclude that $U$ will be added to $\mathcal{L}$ by lines 13-16. Notice that we do not need to explicitly compute $Q_{x}$ as shown in the pseudocode, as we only need the value $\mathrm{vol}(Q_{x})$. Lemma˜5.12 implies that this value can be computed and returned by Algorithm˜1, so we do not need to handle it here.
To see that every set in $\mathcal{L}$ is a $k$-shredder, notice that we only append to $\mathcal{L}$ on line 16. When we append a set $U$ to $\mathcal{L}$, we must have found two components $Q_{x},Q_{y}$ of $G\setminus U$ such that $\mathrm{vol}(Q_{x})+\mathrm{vol}(Q_{y})+|E(U,U)|<m$. By Lemma˜6.8 we have that $U$ is a $k$-shredder, which implies that every set in $\mathcal{L}$ is a $k$-shredder. ∎
@@
Algorithm˜3 runs in $\mathcal{O}(k^{2}m\log^{2}n)$ time.

### v2

@@ 5.4 Additional Lemmas
Suppose that Algorithm 1 on input $(x,\nu,\Pi)$ returns an unverified set $U$ such that $U\neq\varnothing$. Then, $U$ is a $k$-separator. Specifically, let $z$ be the far-most endpoint of an arbitrary path $\pi\in\Pi$. Then, $x$ and $z$ are not connected in $G\setminus U$.
@@
Suppose that Algorithm 1 returns a nonempty unverified set $U$ on input $(x,\nu,\Pi)$. Let $Q_{x}$ denote the component of $G\setminus U$ containing $x$. There exists a modified version of Algorithm 1 that also computes $\mathrm{vol}(Q_{x})$. The modification requires $\mathcal{O}(k\nu)$ additional time, subsumed by the runtime of Algorithm 1.
@@ 6 Resolving Unverified Sets
We now address the case where Algorithm 1 returns an unverified set $U$ on input $(x,\nu,\Pi)$. The difference between $U$ and the list of returned $k$-shredders $\mathcal{L}$ is that the algorithm did not find a component $Q$ of $G\setminus\Pi$ such that $N(Q)=U$. In some sense, this means that we “lose” a component of $G\setminus U$. The danger in guessing that $U$ is a $k$-shredder is that $U$ might only be a $k$-separator. Imagine that $G\setminus U$ has exactly two components: $Q_{1}$ and $Q_{2}$. If $\mathrm{vol}(Q_{2})\gg\mathrm{vol}(Q_{1})$, then it becomes quite challenging to determine whether there exists a third component of $G\setminus U$ in $\mathcal{O}(\mathrm{vol}(Q_{1}))$ time. To cope with this uncertainty, we make another structural classification.
@@
One aspect of this definition may seem strange: we have specifically described $\nu$ as a power of two. This is because in the later sections, we will use geometric sampling to capture $k$-shredders. The sampling parameters we use will be powers of two. Hence, we have imposed this slightly arbitrary detail on Definition 6.1. For now, all that matters is that $\frac{1}{2}\nu<\mathrm{vol}(\mathcal{R})\leq\nu$.
@@
There are two main lemmas for this section. Lemma 6.3 handles low-degree unverified sets and is presented in Section 6.1. Lemma 6.9 handles high-degree unverified sets and is presented in Section 6.2. The idea is that after Algorithm 1 returns an unverified set $U$, we will use Lemma 6.3 to test whether $U$ is a low-degree $k$-shredder. If not, we will leave keep $U$ as a potential high-degree $k$-shredder. After all unverified sets have been returned, we use Algorithm 3 to extract all high-degree $k$-shredders from the remaining unverified sets.
@@ 6.1 Low-Degree
Suppose that $U$ is a low-degree $k$-shredder with partition $(C,\mathcal{R})$. Suppose that the tuple $(x,\nu,\Pi)$ captures $U$ and Algorithm 1 reports $U$ as an unverified set. Our goal is to design an algorithm that confirms $U$ is a $k$-shredder in the same time complexity as Algorithm 1 up to $\text{polylog}(n)$ and $\text{poly}(k)$ factors.
Lemma 5.11 gives us two vertices in different components of $G\setminus U$: $x$ and $z$ (where $z$ is the far-most endpoint of any path in $\Pi$). Since $U$ is a $k$-shredder, there must exist a third component of $G\setminus U$ and every vertex in $U$ must be adjacent to a vertex in this third component. Because $U$ is low-degree, there must exist a vertex $u\in U$ with $\deg(u)\leq\nu$. The idea is to scan through the edges adjacent to this low-degree vertex $u$ and find an edge $(u,y)$ such that $y$ is neither connected to $x$ nor $z$ in $G\setminus U$. To make the scanning procedure viable, we need to efficiently answer pairwise connectivity queries in $G\setminus U$. The key ingredient is a result obtained by Kosinas in [Kos23]. Specifically, the following holds.
@@
We defer the method of capturing $k$-shredders for later. For now, let us assume that the tuple $(x,\nu,\Pi)$ captures a $k$-shredder $U$ and $U$ is reported as an unverified set by Algorithm 1. We show an auxiliary algorithm that will be helpful in listing low-degree $k$-shredders.
@@
After preprocessing the input graph using the data structure from 6.2, there exists a deterministic algorithm that takes in as input $(x,\nu,\Pi,U)$, where $U$ is an unverified set returned by Algorithm 1 on input $(x,\nu,\Pi)$. The algorithm outputs true if $U$ is a $k$-shredder and there exists a vertex $u\in U$ such that $\deg(u)\leq\nu$ in $\mathcal{O}(k^{4}\log n+k\nu)$ time.
@@
In [LS22], Long and Saranurak showed the existence of a pairwise connectivity oracle subject to vertex failures with $\mathcal{O}(k^{2}n^{o(1)})$ update time and $\mathcal{O}(k)$ query time. Our usage of Kosinas’s oracle involves making one update operation with a set of size $k$ and $\mathcal{O}(\nu)$ many connectivity queries. Using Long and Saranurak’s version, we can improve the dependency on $k$ in the time complexity of Algorithm 2 to $\mathcal{O}(k^{2}n^{o(1)}+k\nu)$ time.
@@
As mentioned above, the idea is to scan through the edges adjacent to $u$ in order to find a vertex $y$ such that $y$ is neither connected to $x$ nor $z$ in $G\setminus U$ (see Figure 6). We will be utilizing 6.2 to determine whether a pair of vertices $(x,y)$ are connected in $G\setminus U$. Pseudocode is given in Algorithm 2.
@@
$x$- a sample vertex
1:$\nu$ volume parameter
2:$\Pi$ a set of $k$ openly-disjoint paths starting from $x$
3:$U$ unverified set returned by Algorithm 1 on $(x,\nu,\Pi)$Ensuretrue if (1) and (2) are satisfied, false otherwise
4:ZZ (1) $U$ is a $k$-shredder
5:ZZ (2) there exists a vertex $u\in U$ such that $\deg(u)\leq\nu$Ifall vertices in $U$ have degree greater than $\nu$
[…]
7:$u\leftarrow$ a vertex in $U$ with $\deg(u)\leq\nu$
8:$\pi\leftarrow$ the path in $\Pi$ containing $u$
9:$z\leftarrow$ the far-most endpoint of $\pi$
10:
11:$\Phi\leftarrow$ the pairwise connectivity oracle from 6.2 (already preprocessed)
12:update $\Phi$ on vertex failure set $U$Foreach edge $(u,y)$ adjacent to $u$ such that $y\notin U$
13:$q_{x,y}\leftarrow$ connectivity query between $x$ and $y$ in $G\setminus U$
[…]
@@
Let $(x,\nu,\Pi,U)$ be an input tuple to Algorithm 2. Then, the algorithm returns true if and only if $U$ is a $k$-shredder and there exists a vertex $u\in U$ such that $\deg(u)\leq\nu$.
@@
Suppose that $U$ is a $k$-shredder and that there exists a vertex $u\in U$ such that $\deg(u)\leq\nu$. Let $u$ be the vertex found by line 7. Let $\pi$ be the unique path in $\Pi$ that contains $u$ and let $z$ be the far-most endpoint of $\pi$. These are found by lines 8-9. By Lemma 5.11, we have at least two components of $G\setminus U$ so far: $Q_{x}$ (the component containing $x$) and $Q_{z}$ (the component containing $z$). Since $U$ is a $k$-shredder, there exists a third component $Q_{y}$ of $G\setminus U$. Let $y$ be a vertex in $Q_{y}$ adjacent to $u$. Note that $x,y,z$ are each in distinct components of $G\setminus U$, which implies they are pairwise disconnected in $G\setminus U$. We will find $(u,y)$ by scanning through the edges adjacent to $u$. Prior to calling the algorithm, we will initialize a global pairwise connectivity oracle subject to vertex failures as in 6.2. We update the oracle with respect to $U$ on line 12 so we can answer pairwise connectivity queries in $G\setminus U$. The rest is straightforward. The for loop on line 12 iterates through all edges adjacent to $u$. Hence, the edge $(u,y)$ described above will be found. For each edge $(u,y)$, we perform the two pairwise connectivity queries in $G\setminus U$ by lines 13-14. Lastly, we will discover that $x,y,z$ are all pairwise disconnected in $G\setminus U$ and confirm that $U$ is a $k$-shredder by line 14.
Now suppose that $U$ is not a $k$-shredder or there does not exist a vertex $u\in U$ with $\deg(u)\leq\nu$. If $U$ is not a $k$-shredder, then there are exactly two components in $G\setminus U$: the component containing $x$, and the component containing $z$ (the far-most endpoint of any path in $\Pi$). Therefore, for every edge $(u,y)$ we iterate through such that $u\in U,y\notin U$, either $x$ and $y$ are connected in $G\setminus U$ or $y$ and $z$ are connected in $G\setminus U$. Line 14 guarantees that we will not return true. If there does not exist a vertex $u\in U$ with $\deg(u)\leq\nu$, we can return false immediately. This is done by line 5. ∎
@@
Algorithm 2 on input $(x,\nu,\Pi,U)$ runs in $\mathcal{O}(k^{4}\log n+k\nu)$ time.
@@
Line 5 can be implemented in $\mathcal{O}(k)$ time by iterating over all vertices in $U$ and checking the sizes of their adjacency lists. Lines 7-9 can be done in constant time. The update on line 12 takes $\mathcal{O}(k^{4}\log n)$ time according to 6.2. Iterating over the edges adjacent to $u$ takes $\mathcal{O}(\nu)$ time because $\deg(u)\leq\nu$. For each edge, we make two pairwise connectivity queries, each of which takes $\mathcal{O}(k)$ time according to 6.2. This gives the total time bound of $\mathcal{O}(k)+\mathcal{O}(k^{4}\log n)+\mathcal{O}(k\nu)=\mathcal{O}(k^{4}\log n+k\nu)$. ∎
Proof of Lemma 6.3.
The correctness and time complexity of Algorithm 2 are given by Lemma 6.5 and Lemma 6.6. ∎
@@ 6.2 High-Degree
Now we have a tool to obtain a vertex in $\mathcal{R}$. We are not finished yet, as we need to connect this idea with the unverified sets reported by Algorithm 1. Suppose that $U$ was reported by Algorithm 1 on input $(x,\nu,\Pi)$. We have that $x$ is in some component $Q$ of $G\setminus U$. With Lemma 6.7, our goal is to sample a vertex $y$ in a different component $Q_{y}$ of $G\setminus U$. As described above, we can exploit the fact that $U$ is a high-degree $k$-shredder by exploring low-degree vertices in the neighborhood of $y$ to recover $Q_{y}$. At this point, we have essentially recovered two components of $G\setminus U$: $Q$ and $Q_{y}$. To make further progress, we present a short, intuitive lemma.
@@
There exists a randomized Monte Carlo algorithm that takes as input a list $\mathcal{U}$ of tuples of the form $(x,\nu,\Pi,U)$, where $U$ is an unverified set returned by Algorithm 1 on $(x,\nu,\Pi)$. The algorithm returns a list of $k$-shredders $\mathcal{L}$ that satisfies the following. If $(x,\nu,\Pi,U)\in\mathcal{U}$ is a $k$-tuple such that $U$ is a high-degree $k$-shredder, then $U\in\mathcal{L}$ with probability $1-n^{-100}$. The algorithm runs in $\mathcal{O}(k^{2}m\log^{2}n)$ time.
@@
We summarize the argument made above. Fix a tuple $(x,\nu,\Pi,U)$ in $\mathcal{U}$ such that $U$ is a high-degree $k$-shredder with partition $(C,\mathcal{R})$. Let $\nu$ denote the unique power of two satisfying $\frac{1}{2}\nu<\mathrm{vol}(\mathcal{R})\leq\nu$. Because $U$ is high-degree, we have for all $u\in U$, $\deg(u)>\nu$. Let $Q_{x}$ denote the component of $G\setminus U$ containing $x$. Suppose we have obtained a vertex $y$ in a component $Q_{y}\in\mathcal{R}\setminus\{Q_{x}\}$. Since all vertices in $Q_{y}$ have degree at most $\nu$, the idea is to explore all vertices connected to $y$ that have degree at most $\nu$. Because $N(Q_{y})=U$ and each vertex in $U$ has degree greater than $\nu$, we will compute $Q_{y}$ as the set of explored vertices. After computing $Q_{y}$, we will have obtained two components of $G\setminus U$: $Q_{x}$ and $Q_{y}$. In order to confirm that $U$ is a $k$-shredder, all that is left to do is to make sure that $Q_{x}$ and $Q_{y}$ are not the only components of $G\setminus U$. Lemma 6.8 implies this can be done by verifying that $\mathrm{vol}(Q_{x})+\mathrm{vol}(Q_{y})+|E(U,U)|<m$. Now, the final step is to find such a vertex $y$. We use a random sampling procedure for this task. Consider Algorithm 3. The following two lemmas prove the correctness and time complexity of Algorithm 3.
@@
$\mathcal{U}$- a list of tuples of the form $(x,\nu,\Pi,U)$$\mathcal{L}$- a list of $k$-shredders that satisfies the following:
1:ZZ if $(x,\nu,\Pi,U)\in\mathcal{U}$ is such that $U$ is a high-degree $k$-shredder,
2:ZZ then $U\in\mathcal{L}$ with probability $1-n^{-100}$
[…]
5:independently sample an edge $(y,z)$ uniformly at random
6:$Q_{y}\leftarrow$ set of vertices explored by BFS seeded at $y$:
7:     ignore vertices $v$ such that $\deg(v)>\nu$
[…]
12:return $\mathcal{L}$
[…]
@@
Let $\mathcal{U}$ be the list of tuples given as input to Algorithm 3 and let $\mathcal{L}$ be the output list. Suppose $U$ is a $k$-shredder such that there exists a $k$-tuple $(x,\nu,\Pi,U)\in\mathcal{U}$. Then, $U\in\mathcal{L}$ with probability $1-n^{-100}$. Furthermore, every set in $\mathcal{L}$ is a $k$-shredder.
@@
Fix a tuple $(x,\nu,\Pi,U)$ in $\mathcal{U}$ such that $U$ is a high-degree $k$-shredder with partition $(C,\mathcal{R})$. We will show that $U\in\mathcal{L}$ with high probability. Let $Q_{x}$ denote the component in $\mathcal{R}$ containing $x$. Let $Q_{y}$ denote a component of $\mathcal{R}\setminus\{Q_{x}\}$. We will first show that the algorithm obtains a vertex $y\in Q_{y}$ with high probability. Let $\nu$ denote the unique power of two satisfying $\frac{1}{2}\nu<\mathrm{vol}(Q_{y})\leq\nu$. Lemma 6.7 implies that if we sample $400\frac{m}{\nu}\log n$ edges, we will obtain an edge $(y,z)$ such that $y\in Q_{y}$. Although we do not know the explicit value of $\nu$, all powers of two up to $\frac{m}{\nu}$ are tried by the algorithm during the loop on line 3. Hence, we will eventually obtain an edge $(y,z)$ with $y\in Q_{y}$ with probability $1-n^{-100}$ during the for loop on line 4.
After obtaining a vertex $y\in Q_{y}$, we perform BFS seeded at $y$. Since $\frac{1}{2}\nu<\mathrm{vol}(Q_{y})\leq\nu$, we know that all vertices in $Q_{y}$ have degree at most $\nu$. We also know that $Q_{y}$ can be computed by traversing at most $\nu$ edges because $\mathrm{vol}(Q_{y})\leq\nu$. Lastly, we know that all vertices in $U$ have degree greater than $\nu$ because $U$ is a high-degree $k$-shredder. This implies $Q_{y}$ will be computed by line 6. After computing $Q_{y}$, we can determine whether there exists a tuple $(x,\nu_{x},\Pi,U)$ in the input list such that $N(Q_{y})=U$. If such a tuple exists, we should first check that $x\notin Q_{y}$. If so, we now have two components of $G\setminus U$: $Q_{x}$ and $Q_{y}$. By Lemma 6.8, the only step left to determine whether $U$ is a $k$-shredder is verifying that $\mathrm{vol}(Q_{x})+\mathrm{vol}(Q_{y})+|E(U,U)|<m$. Since $U$ is a $k$-shredder, this inequality must be true. We can conclude that $U$ will be added to $\mathcal{L}$ by lines 9-11. Notice that we do not need to explicitly compute $Q_{x}$ as shown in the pseudocode, as we only need the value $\mathrm{vol}(Q_{x})$. Lemma 5.12 implies that this value can be computed and returned by Algorithm 1, so we do not need to handle it here.
To see that every set in $\mathcal{L}$ is a $k$-shredder, notice that we only append to $\mathcal{L}$ on line 11. When we append a set $U$ to $\mathcal{L}$, we must have found two components $Q_{x},Q_{y}$ of $G\setminus U$ such that $\mathrm{vol}(Q_{x})+\mathrm{vol}(Q_{y})+|E(U,U)|<m$. By Lemma 6.8 we have that $U$ is a $k$-shredder, which implies that every set in $\mathcal{L}$ is a $k$-shredder. ∎
@@
Algorithm 3 runs in $\mathcal{O}(k^{2}m\log^{2}n)$ time.